OpenStudy (anonymous):

Use the pythagorean theorem to determine the length of the unknown side, the perimeter and the area of the triangle. 13 inches, 5 inches, and A

OpenStudy (anonymous):

Are we given the hypotenuse or not?

OpenStudy (anonymous):

Leg 2 Plus Leg 2 = Hypotnuse ^2

OpenStudy (anonymous):

Leg ^2 plus Leg ^2 = Hypotenuse ^2

OpenStudy (anonymous):

yeah but the 13 could be the hypotenuse or a leg. it matters.

OpenStudy (anonymous):

13 is the hypotenuse

OpenStudy (anonymous):

The leg is 5 in

OpenStudy (anonymous):

\[A^2+b^2=c^2\] \[A^2+5^2=13^2\] \[A^2=13^2-5^2\] \[A^2=144\] \[A=12\] Therefore the Area \[=\frac{ 12\times5 }{ 2 }\] \[=30 inches^2\] Perimeter \[=5+12+13\] \[=30inches\]